OpenStudy (anita):

(6x-12)=x+2 --------- 3

OpenStudy (anonymous):

1.multiply 3 on both sides 2.solve for x

OpenStudy (anonymous):

(6x-12)=(x+2)3 6x-12=3x+6 6x-3x=6+12 3x/3=18/3 x=6

OpenStudy (anonymous):

thanks saifoo and hero...=)

OpenStudy (saifoo.khan):

you welcome

OpenStudy (anonymous):

thanks anita...hope i helped you...